From the following information, calculate GDP, private disposable income and private savings:
Consumption 4500
Compensation of employees 6300
Proprietor’s income 900
Corporate profits 700
Net interest 400
Indirect business tax 500
Consumption of fixed capital 1000
Transfers received from government 350
Direct taxes 300
1
Expert's answer
2021-07-22T10:01:45-0400

Using income method, national income (NI) is calculated as follows:

NI=Compensation of employees + Corporate profit +

 Proprietors income + Net interest

"=6300+900+700+400\\=8300"

From national income, GDP is calculated as follows:

GDP =NI + Indirect taxes + Consumption of fixed capital

"=8300+500+1000\\=9800"


Private disposable income=NI +Transfer payment− Direct taxes

"=8300+350-300\\=8350"


Saving =Private disposable income− Consumption

"=8350-45000\\=3850"